OpenStudy (faiqraees):

If the roots of the equation \[\large\rm x^4-px^3+qx^2-pqx+1=0\] are a,b,c and d, show that \[\large\rm (a+b+c)(a+b+d)(b+c+d)(a+c+d)=1\]

OpenStudy (faiqraees):

I dont know if this approach is correct one but this is my view. Let x = 1 (since the equation is true for every value of x) which gives us 1-p+q-pq+1=0 p-q+pq=2 Where p = Σa q= Σab And also pq=Σabc (I dont think this equation would help)

ganeshie8 (ganeshie8):

The given equation is not true for every value of x. It is only true for four values of x.

ganeshie8 (ganeshie8):

namely a, b, c and d.

OpenStudy (welshfella):

right

ganeshie8 (ganeshie8):

Step1 : Notice that a + b + c + d = p. Therefore (a+b+c)(a+b+d)(b+c+d)(a+c+d) = (p-a)(p-b)(p-c)(p-d)

OpenStudy (faiqraees):

@ganeshie8 We have to show the resulting equation. Thus our solution should start from given the equation. Not the way backwards.

jhonyy9 (jhonyy9):

ganeshie8 but we know that than a,b,c and d are roots of this equation so this mean that (x-a)(x-b)(x-c)(x-d)= x^4 -px^3 +qx^2 -pqx +1 too yes ?

ganeshie8 (ganeshie8):

I took only the left hand side. Not the entire equation.

OpenStudy (faiqraees):

@jhonyy9 yes. But here we have too many unnecessary variables

OpenStudy (faiqraees):

@ganeshie8 Okay

ganeshie8 (ganeshie8):

Do you still feel the starting of the proof is wrong ?

jhonyy9 (jhonyy9):

@ganeshie8 possible one bad idea but dont help here using the Viéte formulas ?

OpenStudy (faiqraees):

No its just that, we are not allowed to take in account with the given solution. As if we never heard of given solution

ganeshie8 (ganeshie8):

Yeah, I'm also trying to use Viete formulas and see if we can simplify

jhonyy9 (jhonyy9):

ok. so this mean that make sense really ?

ganeshie8 (ganeshie8):

Are you saying we cannot start the proof with left hand side of the statement that we want to prove ?

OpenStudy (faiqraees):

Yes

ganeshie8 (ganeshie8):

Why not ?

ganeshie8 (ganeshie8):

Do you think it is wrong to do that or you just don't want to do it that way ?

OpenStudy (drmoss):

young man this is not the place to cheat on your test

OpenStudy (faiqraees):

@ganeshie8 Btw I only care about solution. I dont care if it came by bribing a mathematician so you should carry on with your solution. My opinion:- It feels okay. If we prove right to be equal to left it pretty much means the same

OpenStudy (faiqraees):

@DrMoss Not a test. It just I have bounded myself with some rules so I wont be surprised if a type of question came up in which this tactic rendered unusable.

ganeshie8 (ganeshie8):

Okay, so we're good with step1. Lets keep going.

OpenStudy (faiqraees):

OKay

ganeshie8 (ganeshie8):

Step2 : (p-a)(p-b)(p-c)(p-d) is a polynomial in "p" whose roots are a, b, c, d. So we can get its expansion simply by replacing "x" by "p" in the given polynomial : (p-a)(p-b)(p-c)(p-d) = p^4-p(p)^3+q(p)^2-pq(p)+1

ganeshie8 (ganeshie8):

simplify and you're done!

OpenStudy (faiqraees):

By simplification the left side is \[\large\rm p^4-(Σa) p^3+(Σab) p^2-(Σabc)p+abcd\ \]

ganeshie8 (ganeshie8):

I suggest you go through steps 1 and 2 again.

OpenStudy (faiqraees):

Which side are we suppose to simplify? Because I dont think any side needs simplification

ganeshie8 (ganeshie8):

(p-a)(p-b)(p-c)(p-d) = p^4-p(p)^3+q(p)^2-pq(p)+1 = ?

OpenStudy (mathmate):

@FaiqRaees It would be nice if we are given the exact and complete wording of the question. A lot of energy would be wasted if people suggest a solution to be told that it is "not allowed". Math is all about precision, so please give all the parameters of the question to have a level playing field.

OpenStudy (faiqraees):

@ganeshie8 How is p=a+b+c+d = (p-a)(p-b)(p-c)(p-d)?

ganeshie8 (ganeshie8):

I never said p = (p-a)(p-b)(p-c)(p-d)

ganeshie8 (ganeshie8):

a + b + c + d = p gives you a + b + c = p - d b + c + d = p - a c + d + a = p - b d + a + b = p - c

OpenStudy (faiqraees):

@mathmate The question wordings are that exactly. It was just that my teacher didn't accepted solution in which instead of simplifying a given equation, the proof was given by expanding the given solution

OpenStudy (faiqraees):

Oh okay that clears a lot

OpenStudy (faiqraees):

So now if I can show that (p-a)(p-b)(p-c)(p-d) = p^4-p(p)^3+q(p)^2-pq(p)+1 my work is done right?

ganeshie8 (ganeshie8):

Ohkay, I see... I should have added more steps

ganeshie8 (ganeshie8):

I have already shown that for you in step2.

ganeshie8 (ganeshie8):

Let me explain step2 again.

OpenStudy (faiqraees):

Yeah that would be very helpful

ganeshie8 (ganeshie8):

We're given that the roots of the polynomial x^4-px^3+qx^2-pqx+1 are a, b, c, d. So we have x^4-px^3+qx^2-pqx+1 = (x-a)(x-b)(x-c)(x-d)

ganeshie8 (ganeshie8):

With me ?

OpenStudy (faiqraees):

Oh got it

OpenStudy (faiqraees):

Replace x with p

ganeshie8 (ganeshie8):

Yep, simple once you see it :)

OpenStudy (faiqraees):

And since we have already proven that those p brackets are just a simplified version of the given solution, the equation equals the solution

OpenStudy (faiqraees):

But our proof isn't complete yet I suppose. We have to prove that the given solution is equal to 1. We haven't done that or did we?

ganeshie8 (ganeshie8):

(p-a)(p-b)(p-c)(p-d) = p^4-p(p)^3+q(p)^2-pq(p)+1 = ?

OpenStudy (faiqraees):

0

ganeshie8 (ganeshie8):

seriously you should have some rest -.-

OpenStudy (faiqraees):

That is given in the equation x^4-px^3+qx^2-pqx+1=0

ganeshie8 (ganeshie8):

Again, that equation is NOT an identity. It is not true for every value of x. It is true only when x = a, b, c, d.

OpenStudy (faiqraees):

Oh yeah, right, I keep forgetting that. But if we input p, then it should output 1. Yeah okay understood thanks

ganeshie8 (ganeshie8):

Why ?

OpenStudy (faiqraees):

Algebraic manipulation(I loathe myself for missing such thing)

ganeshie8 (ganeshie8):

Happens haha. I'm just making sure you got it right :)

OpenStudy (faiqraees):

@ganeshie8 Yeah thanks

jhonyy9 (jhonyy9):

(x-a)(x-b)(x-c)(x-d)= x^4 -px^3 +qx^2 -pqx +1 = x^3(x-p)+qx(x-p) +1 = = (x^3 +qx)(x-p) +1 this dont help anything ?

OpenStudy (faiqraees):

@ganeshie8 I have a confusion for which the book is giving different answer and the people at math.stackexchange are giving different answer. Can I confirm it with you if you have time?

OpenStudy (faiqraees):

\(\color{#0cbb34}{\text{Originally Posted by}}\) @jhonyy9 (x-a)(x-b)(x-c)(x-d)= x^4 -px^3 +qx^2 -pqx +1 = x^3(x-p)+qx(x-p) +1 = = (x^3 +qx)(x-p) +1 this dont help anything ? \(\color{#0cbb34}{\text{End of Quote}}\) I suppose unless there is an x in the expression, it cant help anything. That's maybe why @ganeshie8 first took steps to eliminate x out of the equations

jhonyy9 (jhonyy9):

@FaiqRaees are you on stackexchange too ?

ganeshie8 (ganeshie8):

I'll try, ask

OpenStudy (faiqraees):

@jhonyy9 Yeah because the good members here come online very late at night(according to my timezone)

jhonyy9 (jhonyy9):

how ? are you from Europe same with me ?

OpenStudy (faiqraees):

@ganeshie8 What is the general formula for the coefficient of x^n-2 of a polynomial f(x) having the degree of n?

OpenStudy (faiqraees):

@jhonyy9 Asia

jhonyy9 (jhonyy9):

ok i understand now

ganeshie8 (ganeshie8):

Do you mean the coefficient of x^(n-2) ?

OpenStudy (faiqraees):

Yes

ganeshie8 (ganeshie8):

What things do we know ?

ganeshie8 (ganeshie8):

Are we given the roots of the polynomial ?

OpenStudy (faiqraees):

No I am just asking generally like in the form Σab or in that way

ganeshie8 (ganeshie8):

You must ask the question clearly. Then only we can pursue the answer...

ganeshie8 (ganeshie8):

If we knew the roots, we could simply use the Vieta formulas to find the coefficients.

OpenStudy (faiqraees):

I dont really know how to describe it Like for an equation x^n +ax^n-1+bx^n-2...... The coefficient of x^(n-1)is Σa What is for x^(n-2)?

ganeshie8 (ganeshie8):

Okay, I see what you're asking.

ganeshie8 (ganeshie8):

Consider the polynomial \[P(x) = x^n + a_{n-1} x^{n-1} + a_{n-2}x^{n-2} + \cdots + a_1 x^1 + a_0 \] Let \(\beta_1, \beta_2, \ldots , \beta_n\) be the roots of \(P(x)\).

OpenStudy (faiqraees):

Okay

ganeshie8 (ganeshie8):

Then we can factor P(x) as \[x^n + a_{n-1} x^{n-1} + a_{n-2}x^{n-2} + \cdots + a_1 x^1 + a_0 = (x-\beta_1)(x - \beta_2)\cdots (x-\beta_n) \]

OpenStudy (faiqraees):

True

ganeshie8 (ganeshie8):

What is the coefficient of x^n on left hand side ?

OpenStudy (faiqraees):

1

ganeshie8 (ganeshie8):

What is the coefficient of x^n on right hand side ?

OpenStudy (faiqraees):

1

ganeshie8 (ganeshie8):

They are equal. So we're good with the coefficients of x^n. Let's compare the coefficients of x^(n-1) both sides.

ganeshie8 (ganeshie8):

What is the coefficient of x^(n-1) on left hand side ?

OpenStudy (faiqraees):

a(n-1)

ganeshie8 (ganeshie8):

What is the coefficient of x^(n-1) on right hand side ?

OpenStudy (faiqraees):

it should be a(n-1) (considering we are allowed to assume the equality to be true otherwise no way to find out, at least for me)

ganeshie8 (ganeshie8):

Right hand side, you should get an expression in terms of the roots. Lets see

ganeshie8 (ganeshie8):

Our goal is to find the coefficient of x^(n-1) in \[ (x-\beta_1)(x - \beta_2)\cdots (x-\beta_n) \]

ganeshie8 (ganeshie8):

How many factors are there in that product ?

OpenStudy (faiqraees):

n

ganeshie8 (ganeshie8):

Yes. How did we find the coefficient of x^n ?

OpenStudy (faiqraees):

For right side? By simple logic. There is no coefficient with x in any of the factors. To reach x^n, the x terms can only multiply with x terms. Since there is no coefficient on any value of x thus the coefficient of x^n is 1. Or you can also say that by just comparing it with left side